Circular No. 7345

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

MARKARIAN 421 C. Gouiffes, Centre d’Etudes de Saclay, Gif-sur-Yvette, forwards the

following message for B. Degrange, who writes on behalf of the CAT collaboration (Centre d’Etude Spatiale des Rayonnements, Toulouse; Collège de France, Paris; Laboratoire de Physique Nucléaire et des Hautes Energies (LPNHE) Ecole Polytechnique; LPNHE, Universités Paris VI et VII; Service d’Astrophysique, Centre d’Etudes de Saclay; University of Perpignan; Charles University, Prague; and Olomouc University): “High-TeV activity has been observed recently from the BL Lac object Mkn 421, using the very-high-energy γ-ray CAT telescope located in the French Pyrenées and operating at an energy threshold of 250 GeV. Mkn 421 was observed on Jan. 7, 11, 12, and 13 with a total on-source exposure time of 12.5 hr. A preliminary analysis indicates that the source was variable on a daily timescale with the following average nightly flux levels, in units of the Crab Nebula flux: Jan. 7.13 UT, 0.45; 11.15, 2.3; 12.1, 1.9; 13.1, 1.3.”

SUPERNOVA 1999gs IN NGC 4725 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in NGC 4725. The new object was discovered at mag ∼ 19.3 on an unfiltered image (limiting mag ∼ 19.5) taken on 1999 Dec. 28.5 UT, and it was confirmed on a 5-min unfiltered image (limiting mag ∼ 21.0) taken on Dec. 30.5. SN 1999gs is located in the southern spiral arm of the galaxy at α = 12^{h}50^{m}26^{s}.28, δ = +25^{o}28′16′′.5 (equinox 2000.0), which is 3′′ west and 105′′ south of the nucleus of NGC 4725. Another observation obtained on 2000 Jan. 13.5 showed the new object still at mag ∼ 19.5 (limiting mag 20.5). A template image of NGC 4725, made by combining several observations of the field before 1999 June 7, showed nothing at the position of SN 1999gs (limiting mag ∼ 20.5). The discovery images can be found

at http://astron.berkeley.edu/~bait/1999/ngc4725.html. LOSS has

not observed this galaxy for about half a year, so SN 1999gs must be a very old supernova.

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates by A. Pearce, Nedlands, W. Australia (0.41-m

reflector): 1999 Nov. 1.80 UT, 12.9; Dec. 14.81, 12.9; 2000 Jan. 8.70, 13.3.

2000 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green